# Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth

### v1

@@ 3.1 Objectives and Problem Statement
In practice, the principal may collect responses from multiple agents, resulting in an $n\times m$ response matrix $\hat{X}$ where each entry $\hat{X}_{i,q}\in\Sigma$ if agent $i$ answers question $q$ and $\hat{X}_{i,q}=\phi$ otherwise. Given $\hat{X}$ and the principal’s samples of the cheap signal $Z$, a scoring mechanism designs a score for each agent. Our goal is to design a scoring mechanism that satisfies information monotonicity. Let $S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ denote the expected score of agent $i$, where $\hat{X}_{-i}$ is the reports of all agents but $i$.
@@
A mechanism is information monotone for naive strategies if, for any agent $i$, $S_{i}(X_{i},X_{-i}\mid Z)>S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ for any naive strategy $\nu_{i}\in\Theta_{\nu}$ with $\hat{X}_{i}=\nu_{i}(X_{i},Z_{i})$.
A mechanism is $\epsilon$-information monotone if it is information monotone for naive strategies and, additionally, $S_{i}(X_{i},X_{-i}\mid Z)\geq S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)-\epsilon$ holds for any strategy $\theta_{i}\in\Theta$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$.
At the individual level, information monotone scores can be used to identify unreliable labelers or LLM cheaters. At the dataset level, the score distribution across agents can help interpret the quality of the crowdsourced dataset.
We note that the same concept is termed informed truthfulness or informed strategy-proofness in the peer prediction literature, where the objective is to incentivize agents to report truthful information (Shnayder et al., 2016).
@@ 4 Our Method
Classic peer prediction methods often struggle with cheap signals that are highly correlated with one another but only weakly correlated with the ground truth. In this section, we generalize a prior work — the correlated agreement (CA) mechanism (introduced in Appendix A) — and propose a mechanism that achieves information monotonicity when the principal can obtain some noisy samples of the cheap signal. In the context of LLM cheating, these samples can be LLM-generated labels for the crowdsourced questions. . We name our generalization as the conditioned CA mechanism, which scores an agent based on the correlation between her responses and her peers’ conditioned on the principal’s samples of $Z$. Our method has two main steps.
Learn the scoring function. A scoring function determines whether a pair of signals “agree” or not, which is learned based on data. We first estimate the empirical conditioned joint distribution $\tilde{P}(\hat{X}_{i},\hat{X}_{j}\mid Z)$, computed using the frequency of observing a pair of signals on the same question conditioned on $Z$.***We can use all agents’ reports to estimate the same distribution because we assume agents are homogeneous and questions are i.i.d.. This assumption is primarily made for practical concerns, as there is usually a shortage of data to estimate the distribution accurately for each pair of agents. Next, we compute the delta tensor
@@
which is the difference between the joint distribution and the product of marginal distributions of the reports, conditioned on $Z$. This gives us the scoring function $T_{\tilde{\Delta}}=\text{Sign}(\tilde{\Delta})$, i.e. $T_{\tilde{\Delta}}(h,l,k)=1$ if $\tilde{\Delta}_{h,l,k}>0$ and 00 otherwise. This step becomes unnecessary if the true scoring function $T_{\Delta}=\text{Sign}(\Delta)$ is already known, where $\Delta$ is defined analogously to $\tilde{\Delta}$, but based on the true signals $X_{i}$ and $X_{j}$. For example, for binary questions with positively correlated signals, $\Delta_{\cdot,\cdot,k}$ is an identity matrix for any $k\in\Sigma$.
Compute the bonus/penalty score. We then repeatedly sample a bonus question $q$ and a penalty question $q^{\prime}$ while ensuring $Z_{q}=Z_{q^{\prime}}=k$. For each pair of samples, the score for agent $i$ is:
@@
where $j$ is a randomly selected agent who answers question $q^{\prime}$. The final score for agent $i$ is obtained by averaging over all sampled pairs of $q$ and $q^{\prime}$. Algorithm 1 presents the procedures of our method.
for agent $i\in[n]$ do
       Let $M_{i}$ be the set of questions answered by agent $i$ Initialize $s_{i}=0$;
       for $k\in\Sigma$ do
             Initialize $x=0$;
             for question $q\in M_{i}\cap M_{Z}(k)$ do
                   Randomly select a peer $j$ who answers $q$;
                   $x\mathrel{+}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q},k)$;
                   Randomly select a question $q^{\prime}\in M_{j}\cap M_{Z}(k)$;
                   $x\mathrel{-}=T_{\tilde{\Delta}}(\hat{X}_{i,q},\hat{X}_{j,q^{\prime}},k)$
            $s_{i}\mathrel{+}=x\cdot\tilde{P}(Z=k)/|M_{i}\cap M_{Z}(k)|$
return $s$
Algorithm 1 The conditioned CA mechanism
Intuitively, the conditioned CA mechanism encourages agents to agree on the same question and penalizes agents when they agree on two distinct questions, given that the principal’s sampled signal is identical on both questions. If agents rely purely on the cheap signal or respond independently of the true signal, the bonus score and the penalty score cancel out in expectation. This results in a score that is close to zero.
@@ 4.1 Theoretical Insights
We demonstrate the effectiveness of our method and clarify the conditions under which the mechanism performs reliably. Before we start, we first present an intuitive way to interpret the score computed by our mechanism. In a prior work, Kong and Schoenebeck (2019a) show that the score of the CA mechanism can be interpreted as a type of mutual information between agents’ reports called the total variation distance (TVD) mutual information. Inspired by this prior work, we show that when agents report the high-effort signal and the underlying scoring function $T_{\Delta}$ is applied, the expected score computed by the conditioned CA mechanism is essentially the TVD conditioned mutual information:
@@ 4.1.1 Conditions for $\epsilon$-Information Monotonicity
We assume $Z_{i}$ and $Y$ are approximately independent conditioned on $Z$, i.e. $|\Pr(Z_{i},Y\mid Z)-\Pr(Z_{i}\mid Z)\Pr(Y\mid Z)|\leq\epsilon_{i}$ for any $i\in[n]$.
@@
We defer the proof to Section B.1. When agents report the high-effort signals truthfully, the expected score is the TVD mutual information between $X_{i}$ and $X_{j}$. Theorem 4.3 suggests that if $Z$ approximately blocks the correlation between $Z_{i}$ with $X_{j}$ and $Z_{j}$, then any cheating strategy can only decrease the TVD mutual information between the reports $\hat{X}_{i}$ and $\hat{X}_{j}$ up to an error. In the context of LLM cheating, suppose the principal suspects that some agents are using LLMs to generate their labels. Our mechanism suggests that the principal can generate labels using a (potentially different) LLM and compute the CA score conditioned on these generated labels. According to Theorem 4.3, if the principal’s generated labels satisfy two conditions — 1) they block correlations among the labels generated by potential cheating agents (4.2), and 2) they block correlations between those labels and the ground truth (4.1) — then the conditioned CA mechanism’s score is information monotone.
We further show the necessity of the assumptions for information monotonicity. We construct an example to show that if $Z_{i}$ and $Z_{j}$ have some correlation that is not captured by $Z$, then there exists a signal structure such that the score computed by the conditioned CA mechanism is not maximized by reporting $X_{i}$ and $X_{j}$.
@@
Suppose $T_{\Delta}$ is known. If there exist $(z_{i},z_{j},y,z)$ such that $|\Pr(Z_{i}=z_{i},Z_{j}=z_{j}\mid Y=y,Z=z)-\Pr(Z_{i}=z_{i}\mid Y=y,Z=z)\Pr(Z_{j}=z_{j}\mid Y=y,Z=z)|=d$ for some $0<d\leq\frac{1}{4}$, then there exists a joint distribution $\Pr(X_{i},X_{j},Z_{i},Z_{j},Y,Z)$ that satisfies the causal relationship described in Fig. 1 and a strategy profile $(\theta_{i},\theta_{j})\neq(\tau,\tau)$ such that $S_{i}(\hat{X}_{i},\hat{X}_{j})\geq S_{i}(X_{i},X_{j})+8d^{2}$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ and $\hat{X}_{j}=\theta_{j}(X_{j},Z_{j})$.
Before we provide the full proof (deferred to Section B.2), we briefly introduce the high-level idea. Consider a signal structure with a binary signal space $\Sigma=\{0,1\}$. Suppose agent $i$’s high-effort signal $X_{i}$ is perfectly accurate at predicting $X_{j}$ when $X_{i}=0$ but is noisy when $X_{i}=1$, i.e. $\Pr(X_{j}=0|X_{i}=0,Z)=1$ and $\Pr(X_{j}=1|X_{i}=1,Z)<1$. On the other hand, the cheap signal $Z_{i}$ is perfectly accurate at predicting $Z_{j}$ when $Z_{i}=1$ but is noisy when $Z_{i}=0$, i.e. $\Pr(Z_{j}=1|Z_{i}=1,Z)=1$ and $\Pr(Z_{j}=0|Z_{i}=0,Z)<1$. In this case, it is intuitive that agent $i$ can be better off if both agents combine their signals — reporting 0 if and only if $X_{i}=Z_{i}=0$ and reporting $1$ otherwise. Note that although this example only indicates that 4.2 is necessary, the same recipe can be straightforwardly applied to show the necessity of 4.1.
4.1.2 Conditions for Information Monotonicity Under Naive Strategies
We know 4.1 and 4.2 are necessary for the conditioned CA mechanism to be $\epsilon$-information monotone. However, these assumptions are designed to guard against complex, unrealistic strategies, such as those where cheaters first obtain the high-effort signal and then blend it with the cheap signal. If we instead restrict our focus to naive strategies, these assumptions on $Z$ can be significantly relaxed.
@@ 4.1.2 Conditions for Information Monotonicity Under Naive Strategies
For any pair of agents $i$ and $j$, we assume the TVD conditioned mutual information between $X_{i}$, $X_{j}$, $Z_{i}$, $Z_{j}$ satisfies that
@@
If $T_{\Delta}$ is known and 4.5 holds, the conditioned CA mechanism is information monotone for naive strategies.
We defer the proof to Section B.3. Intuitively, a naive strategy is essentially a mixture of truth-telling and uninformative strategies. We show that when both agents follow naive strategies, the score-maximizing equilibrium must fall into one of the following three cases: 1) both agents exert full effort, achieving an expected score of $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$; 2) both agents report cheap signals, with expected score $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$; 3) one agent exerts full effort while the other reports a cheap signal, yielding expected score $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. Therefore, the inequality in 4.5 ensures that full effort strictly dominates naive strategies, establishing the proposition.
@@
4.1.3 Learning the scoring function $T_{\Delta}$.
Thus far, we have assumed that the true scoring function $T_{\Delta}$ is known. For complex signal structures, especially when the signal space is large, we have to learn the scoring function from the data. This introduces a concern: if a significant fraction of agents are cheating, the learned scoring function $T_{\tilde{\Delta}}$ might be manipulated in a way that benefits the cheating reports. We mitigate this concern using the following proposition. Let $S_{i}^{T}(\theta_{i},\theta_{j})$ be the expected score computed under the scoring function $T$ when agents’ strategies are $\theta_{i}$ and $\theta_{j}$ respectively.
Proposition 4.7.
We defer the proof to Section B.4. Intuitively, the proposition holds because when $T_{\Delta}$ is known and agents are truth-telling, the expected score is $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$. However, if $T_{\Delta}$ is mislearned, the expected score is upper bounded by $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)$. Since our assumptions ensure that $Z_{i}$ and $Z_{j}$ are independent given $Z$, the data processing inequality guarantees that $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j}\mid Z)<I_{\text{TVD}}(X_{i};X_{j}\mid Z)$.
4.7 implies that the expected score of the conditioned CA mechanism is maximized if every agent exerts full effort and reports $X_{i}$, while any strategy that manipulates the true scoring function will only (weakly) decrease the expected score.***Here we assume that the principal can learn the correct $T_{\Delta}$ if everyone reports their high-effort signal.
@@ 5 Experiments
We evaluate the effectiveness of our proposed method using real-world crowdsourcing datasets. We focus on labeling tasks where human judgment remains as the gold standard, given that current LLMs still fall short of perfect performance. We first empirically examine the conditions under which the key assumptions underlying our method are likely to hold, as well as when they may fail. Next, we simulate a mixed, noisy crowd of labelers and assess whether our method can detect unreliable labelers better than several well-established baselines.
@@ 5.1 Datasets
Hatefulness/Offensiveness/Toxicity Labeling  We first consider a toxicity labeling dataset which includes $3459$ social media user comments posted in response to political news posts and videos on Twitter, YouTube, and Reddit in August 2021 (Schöpke-Gonzalez et al., 2025). MTurk workers are assigned these comments and are requested to annotate them as hateful, offensive, and/or toxic. The definitions of hatefulness, offensiveness, and toxicity are selected from highly cited previous literature. In this dataset, every question is answered by $5$ workers, and each worker answers $73$ questions on average.
Preference Alignment
We further consider an LLM alignment dataset where each question is a comparison between two answers to the same prompt generated by two different LLMs. Labelers rate their preference on a 0–4 scale (0: A clearly better, 4: B clearly better, 2: tie). The comparison is scored from several more detailed angles, such as the helpfulness and the truthfulness of the answer. Nonetheless, we only focus on the overall preference. In total, there are 10 4611046110\,46110 461 questions where each question is answered by two normal labelers and two experts. On average, each labeler answers around $180$ questions. All labelers are recruited from Prolific. For more details, refer to Miranda et al. (2024).
5.2 Large Language Model Usage
In this work, we select five well-known LLMs to generate labels or annotations for all used datasets, including GPT-3.5-turbo, GPT-4, Gemma-2-2b-it***https://huggingface.co/google/gemma-2-2b-it, Phi-3.5-mini-Instruct ***https://huggingface.co/microsoft/Phi-3.5-mini-instruct, and Mistral-7B-Instruct-v0.3***https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3. The corresponding prompt templates are provided in Appendix E. Note that we choose only five models because other commonly used LLMs, such as LLaMA and Gemini, refuse to answer toxicity-related questions. In Appendix C, we use GPT-4 to process each human review and LLM-generated review into a $19$-dimension vector to implement our method.

### v2

@@ 3.1 Objectives and Problem Statement
In practice, the principal may collect responses from multiple agents, resulting in an $n\times m$ response matrix $\hat{X}$ where each entry $\hat{X}_{i,k}\in\Sigma$ if agent $i$ answers question $k$ and $\hat{X}_{i,k}=\emptyset$ otherwise. Given $\hat{X}$ and the principal’s samples of the cheap signal $Z$, a scoring mechanism designs a score for each agent. Our goal is to design a scoring mechanism that satisfies information monotonicity. Let $S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ denote the expected score of agent $i$, where $\hat{X}_{-i}$ is the reports of all agents but $i$.
@@
A mechanism is information monotone for lazy-reporting strategies if, for any agent $i$, $S_{i}(X_{i},X_{-i}\mid Z)>S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)$ for any lazy-reporting strategy $\nu_{i}\in\Theta_{\nu}$ with $\hat{X}_{i}=\nu_{i}(X_{i},Z_{i})$.
A mechanism is $\epsilon$-information monotone if it is information monotone for lazy-reporting strategies and, additionally, $S_{i}(X_{i},X_{-i}\mid Z)\geq S_{i}(\hat{X}_{i},\hat{X}_{-i}\mid Z)-\epsilon$ holds for any strategy $\theta_{i}\in\Theta$ with $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$.
At the individual level, information monotone scores can be used to identify low-effort agents. At the dataset level, the score distribution across agents can help interpret the quality of the crowdsourced dataset.
We note that the same concept is termed informed truthfulness or informed strategy-proofness in the peer prediction literature, where the objective is to incentivize agents to report truthful information shnayder2016informed.
@@ 4 Our Method
Classic peer prediction methods often struggle with cheap signals, which are highly correlated with one another but only weakly correlated with the ground truth. In this section, we generalize a prior work—the correlated agreement (CA) mechanism (introduced in Appendix B)—and propose a mechanism that achieves information monotonicity when the principal can obtain some noisy samples of the cheap signal. In the context of LLM contamination, these samples can be the labels generated by a popular LLM on a fraction of the crowdsourced questions.
We call our generalization the conditioned CA mechanism, which scores an agent based on the correlation between her responses and her peers’ conditioned on the principal’s samples of $Z$. Our method has two main steps.
Learn the scoring function. A scoring function determines whether a pair of signals “agree” or not, which is learned based on data. We first estimate the empirical conditioned joint distribution $\tilde{P}(\hat{X}_{i},\hat{X}_{j}\mid Z)$, computed using the frequency of observing a pair of signals on the same question conditioned on ZZ. [^1] We can use all agents’ reports to estimate the same distribution because we assume agents are homogeneous and questions are i.i.d.. This assumption is primarily made for practical concerns, as there is usually a shortage of data to estimate the distribution accurately for each pair of agents. Next, we compute the delta tensor
@@
[…]
Compute the bonus/penalty score. Fixing a $k\in\Sigma$, we then repeatedly sample a bonus question $q$ and a penalty question $q^{\prime}$ while ensuring $Z_{q}=Z_{q^{\prime}}=k$. For each pair of samples, the score for agent $i$ is:
@@
where $j$ is a randomly selected agent who answers both $q$ and $q^{\prime}$. The final score for agent $i$ is obtained by averaging over all $k\in\Sigma$ and the sampled pairs of $q$ and $q^{\prime}$.
Intuitively, the conditioned CA mechanism encourages agents to agree on the same question and penalizes agents when they agree on two distinct questions, given that the principal’s sampled signal is identical on both questions. If agents rely purely on the cheap signal or respond independently of the true signal, the bonus score and the penalty score cancel out in expectation. This results in a score that is close to zero. We defer more details to Appendix C.
@@ 4.1 Theoretical Insights
We demonstrate the effectiveness of our method and clarify the conditions under which the mechanism performs reliably. Before we start, we first present an intuitive way to interpret the score computed by our mechanism. In a prior work, FrameworkKong2019 show that the score of the CA mechanism can be interpreted as a type of mutual information between agents’ reports called the total variation distance (TVD) mutual information. Inspired by this prior work, we show that when agents report the high-effort signal and the underlying scoring function $T_{\Delta}$ is applied, the expected score computed by the conditioned CA mechanism is essentially the conditional TVD mutual information:
@@ 4.1.1 Conditions for $\epsilon$-Information Monotonicity
We assume $Z_{i}$ and $Y$ are approximately independent conditioned on $Z$, i.e. $\left|\Pr(Z_{i},Y\mid Z)-\Pr(Z_{i}\mid Z)\Pr(Y\mid Z)\right|\leq\epsilon_{i}$ for any $i\in[n]$.
@@
We defer the proof to Section D.1. When agents report the high-effort signals truthfully, the expected score is the TVD mutual information between $X_{i}$ and $X_{j}$. Theorem 4.3 suggests that if $Z$ approximately blocks the correlation between $Z_{i}$ with $X_{j}$ and $Z_{j}$, then any manipulation can only decrease the TVD mutual information between $\hat{X}_{i}$ and $\hat{X}_{j}$ up to an error.
We further show the necessity of the assumptions for information monotonicity. The following proposition suggests that if $Z_{i}$ and $Z_{j}$ have some correlation that is not captured by $Z$, then there exists a signal structure following the relationship in Figure 1, such that the score computed by the conditioned CA mechanism is not maximized by reporting $X_{i}$ and $X_{j}$.
@@
Suppose $T_{\Delta}$ is known. If there exist $(z_{i},z_{j},y,z)$ such that $|\Pr(Z_{i}=z_{i},Z_{j}=z_{j}\mid Y=y,Z=z)-\Pr(Z_{i}=z_{i}\mid Y=y,Z=z)\Pr(Z_{j}=z_{j}\mid Y=y,Z=z)|=d$ for some $0<d\leq\frac{1}{4}$, then there exists a joint distribution $\Pr(X_{i},X_{j},Z_{i},Z_{j},Y,Z)$ that satisfies the causal relationship described in Figure 1 and a strategy profile $(\theta_{i},\theta_{j})\neq(\tau,\tau)$ such that $S_{i}(\hat{X}_{i},\hat{X}_{j}\mid Z)\geq S_{i}(X_{i},X_{j}\mid Z)+8d^{2}$ where $\hat{X}_{i}=\theta_{i}(X_{i},Z_{i})$ and $\hat{X}_{j}=\theta_{j}(X_{j},Z_{j})$.
Before we provide the full proof (deferred to Section D.2), we briefly introduce the high-level idea. Consider a signal structure with a binary signal space $\Sigma=\{0,1\}$. Suppose agent $i$’s high-effort signal $X_{i}$ is perfectly accurate at predicting $X_{j}$ when $X_{i}=0$ but is noisy when $X_{i}=1$, i.e. $\Pr(X_{j}=0|X_{i}=0,Z)=1$ and $\Pr(X_{j}=1|X_{i}=1,Z)<1$. On the other hand, the cheap signal $Z_{i}$ is perfectly accurate at predicting $Z_{j}$ when $Z_{i}=1$ but is noisy when $Z_{i}=0$, i.e. $\Pr(Z_{j}=1|Z_{i}=1,Z)=1$ and $\Pr(Z_{j}=0|Z_{i}=0,Z)<1$. In this case, it is intuitive that agent $i$ can be better off if both agents combine their signals—reporting 0 if and only if $X_{i}=Z_{i}=0$ and reporting $1$ otherwise. Note that although this example only indicates that 4.2 is necessary, the same recipe can be straightforwardly applied to show the necessity of 4.1.
4.1.2 Conditions for Information Monotonicity Under Lazy-reporting Strategies
We know 4.1 and 4.2 are necessary for the conditioned CA mechanism to be $\epsilon$-information monotone. However, these assumptions are designed to guard against complex, unrealistic strategies, such as those where agents first obtain the full-effort signal and then blend it with the cheap signal. If we instead restrict our focus to lazy-reporting strategies, these assumptions on $Z$ can be significantly relaxed.
@@ 4.1.2 Conditions for Information Monotonicity Under Lazy-reporting Strategies
For any pair of agents $i$ and $j$, we assume the conditional TVD mutual information between $X_{i}$, $X_{j}$, $Z_{i}$, $Z_{j}$ satisfies that
@@
If $T_{\Delta}$ is known and 4.5 holds, the conditioned CA mechanism is information monotone for lazy-reporting strategies.
We defer the proof to Section D.3. Intuitively, a lazy-reporting strategy is essentially a mixture of truth-telling and no-effort strategies. We show that when both agents follow lazy-reporting strategies, the score-maximizing equilibrium must fall into one of the following three cases: 1) both agents exert full effort, achieving an expected score of $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$; 2) both agents report cheap signals, with expected score $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$; 3) one agent exerts full effort while the other reports a cheap signal, yielding expected score $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. Therefore, the inequality in 4.5 ensures that full effort strictly dominates lazy-reporting strategies, establishing the proposition.
@@
Remark 4.7.
Thus far, we have assumed that the true scoring function $T_{\Delta}$ is known. In Appendix E, we show that the theoretical guarantees still hold even when we have to learn the scoring function from the noisy and potentially corrupted data.
@@ 5 Experiments
We evaluate the effectiveness of our proposed method using real-world crowdsourcing datasets. We focus on labeling tasks where human judgment remains as the gold standard, given that current LLMs still fall short of perfect performance. We first empirically examine the conditions under which the key assumptions underlying our method are likely to hold, as well as when they may fail. Next, we simulate a mixed, noisy crowd of labelers and assess whether our method can detect unreliable labelers better than several well-established baselines. The datasets and code of our experiments are available at https://github.com/yichiz97/LLM_contamination.
@@ 5.1 Datasets
Hatefulness/Offensiveness/Toxicity Labeling  We first consider a toxicity labeling dataset which includes $3459$ social media user comments posted in response to political news posts and videos on Twitter, YouTube, and Reddit in August 2021 (schopke2025using). MTurk workers are assigned these comments and are requested to annotate them as hateful, offensive, and/or toxic. The definitions of hatefulness, offensiveness, and toxicity are selected from highly cited previous literature. Every question is answered by $5$ workers and each worker answers $73$ questions on average.
Preference Alignment  We further use an alignment dataset where each question compares two LLM-generated answers to the same prompt. Labelers rate their preference on a 0–4 scale (0: A clearly better, 4: B clearly better, 2: tie). The comparison is scored from several more detailed angles, while we only focus on the overall preference. The dataset contains 10 46110\,461 questions, each labeled by two crowd workers and two experts. On average, each labeler answers about 180 questions. For more details, refer to miranda2024hybrid.